OpenStudy (anonymous):

Find the sum of n terms of series: 1/2 + 3/4 + 7/8 + 15/16 + ................

OpenStudy (anonymous):

do you know the sum of n terms of series 1/2+1/4+1/8+1/16.....? if so, how is that any different from 1-1/2+1-1/4+1-1/8...

OpenStudy (experimentx):

\[ \sum_{n=1}^\infty {2^n-1 \over 2^n}\] i guess you can do this

OpenStudy (anonymous):

I want to solve it in AGP

OpenStudy (anonymous):

i took 1/2 =x but i am stuck

OpenStudy (experimentx):

diverges!!

OpenStudy (experimentx):

the former part diverges ... so the whole series diverges

OpenStudy (anonymous):

not exactly sure what AGP is but look at 1+1+1+1... for n terms then look at 1/2+(1/2)^2+(1/2)^3+... for n terms then 1-1/2= 1/2 1-(1/2)^2=1/4 ... does this help?

OpenStudy (anonymous):

1-(1/2)^2=3/4*

OpenStudy (experimentx):

is that up to n terms??

OpenStudy (experimentx):

n - 1/2(1/(1 - 1/2)) = n - 1